OpenStudy (anonymous):

Find the vertex of the parabola. State whether the vertex is a max or a min. f(x) = -2x2 + 12x + 13

OpenStudy (mimi_x3):

a = -2 b = 12 c = 13 Substitute it into the formula, then there's your vertex V = (-b/2a , a(-b/2a)^2 + b(-b/2a) +c)

OpenStudy (anonymous):

The vertex of a parabola is the point where the derivative is zero: \[f'(x) = -4x + 12 = 0\] \[x = 3\] that's the x coordinate of our vertex point to find the y coordinate, we must find f(3) (the value of y at x = 3) \[f(3) = -18 + 36 + 13 = 31\] so the vertex of f(x) is at (3,31) to find out whether it is a minimum or a maximum, we find the second derivative of f(x):\[f''(x) = -4\]since f''(x) is negative, the graph is concave down at that poin. Therefore, the vertex is a maximum.

OpenStudy (mimi_x3):

What ? It says find the vertex not the Point Of Inflexion ?

OpenStudy (anonymous):

exactly

OpenStudy (mimi_x3):

So am I wrong ?

OpenStudy (anonymous):

actually, your equation does the exact same thing as my working out. You are also right.

OpenStudy (mimi_x3):

Oh, so you took extra steps xD

OpenStudy (anonymous):

no extra steps. the only extra thing I did was to find out whether the point was a minimum or a maximum, which you didn't do in your answer

OpenStudy (anonymous):

f(x)=-2(x^2-6x)+13 f(x)=-2(x-3)^2+13-(9.(-2)) f(x)=-2(x^2-6x)+13+18 f(x)=-2(x^2-6x)+31 vertex ( 3,31) maximum , open down Domain: all real number \[Range: (-\infty,31)\]

OpenStudy (anonymous):

ty

OpenStudy (mimi_x3):

Yeah, I didn't do the minimum or maximum because I just showed how to do it xD . The person can find the minimum or maximum easily anyway xD